A mass of 0.750 kg is hung from a vertical spring with a spring constant of 20 N/m. After coming to equilibrium the mass is then pulled down an additional distance 10 cm and then released from rest. The mass then oscillates from position 3 to 4 to 5 to 4 and back to 3 over and over.




A) List the types of nonzero energy that the system has at xmin (Position 5). B) Calculate xeq.




C) Calculate hmax.




D) Calculate the total energy of the system at xmax (Position 3).




E) Calculate the maximum speed.




F) Calculate the amounts of EK, Eg, and ES when the mass is 4.00 cm above the equilibrium position.




G) Explain two ways that the maximum speed can be decreased


1
Expert's answer
2022-04-11T12:05:21-0400

"h_{max}=2\u2206x=20~cm,"

"E=\frac{k∆x^2}2=0.1~J,"

"\omega= \sqrt\frac 
km

 =5.2~\frac{rad}s,"

"v_{max}=\omega∆x=0.52~\frac ms ."